# Enhanced DACER Algorithm with High Diffusion Efficiency

### v1

@@ Abstract
Due to their expressive capacity, diffusion models have shown great promise in offline RL and imitation learning. Diffusion Actor-Critic with Entropy Regulator (DACER) extended this capability to online RL by using the reverse diffusion process as a policy approximator, trained end-to-end with policy gradient methods, achieving strong performance. However, this comes at the cost of requiring many diffusion steps, which significantly hampers training efficiency, while directly reducing the steps leads to noticeable performance degradation. Critically, the lack of inference efficiency becomes a significant bottleneck for applying diffusion policies in real-time online RL settings. To improve training and inference efficiency while maintaining or even enhancing performance, we propose a Q-gradient field objective as an auxiliary optimization target to guide the denoising process at each diffusion step. Nonetheless, we observe that the independence of the Q-gradient field from the diffusion time step negatively impacts the performance of the diffusion policy. To address this, we introduce a temporal weighting mechanism that enables the model to efficiently eliminate large-scale noise in the early stages and refine actions in the later stages. Experimental results on MuJoCo benchmarks and several multimodal tasks demonstrate that the DACER2 algorithm achieves state-of-the-art performance in most MuJoCo control tasks with only five diffusion steps, while also exhibiting stronger multimodality compared to DACER.
@@ 1 Introduction
Figure 1: Diffusion steps and performance. Taking the Walker2d-v3 task as an example.
Diffusion models have recently demonstrated remarkable performance in image generation epstein2023diffusion ; zhang2023text ; miao2024training and video generation ceylan2023pix2video ; yang2024cogvideox ; wang2025lavie . Owing to their expressive capabilities, diffusion models can represent a wide range of complex distributions, rendering them highly effective for diverse generative tasks. This expressiveness also positions them as a suitable policy class for continuous control, commonly referred to as diffusion policies ren2024DPPO ; li2024learning ; lu2025makes . Moreover, their generative nature aligns naturally with offline reinforcement learning (RL) paradigms wang2022diffusion ; ada2024diffusion ; kang2023efficient .
@@
In online RL, adopting an energy-based model as the policy equips the agent with powerful representational capabilities. Learning a policy to approximate the corresponding energy-based target distribution allows for modeling complex and multimodal action patterns without relying on restrictive parametric assumptions, especially in continuous action spaces. This enhanced expressiveness can significantly improve exploration by enabling the agent to discover and exploit diverse behavioral strategies. However, effectively approximating such an expressive soft policy presents notable challenges. A key difficulty lies in how to efficiently and accurately sample from the target distribution. While algorithms such as Soft Actor-Critic (SAC) haarnoja2018soft and Distributional Soft Actor-Critic (DSAC) duan2021distributional ; duan2023dsac aim to approximate the soft-target distribution, they generally represent the policy as a simple Gaussian. This choice is computationally efficient but falls short in modeling complex, multimodal behavior.
@@
Recent advances have explored more expressive sampling approaches, such as diffusion models psenka2023learning ; ma2025soft ; messaoud2024s , to better approximate the Boltzmann distribution. Existing methods can be broadly categorized into two groups: score matching and policy gradient approaches. In the first group, QVPO ding2024diffusion proposes using Q-weighted imitation learning samples to improve policy learning. QSM psenka2023learning directly aligns the score functions with the gradients of the learned Q functions and uses Langevin dynamics for sampling. DIPO yang2023diffusion updates the replay buffer using action gradients and improves the performance of the policy through a diffusion loss. In the second group, DACER wang2024diffusion directly backward the gradient through the reverse diffusion process and proposes a GMM entropy regulator to balance exploration and exploitation. However, for online RL agents that require continuous environment interaction, the computational efficiency of these methods during inference becomes a critical concern. Score-matching-based methods typically require 20 diffusion steps, sometimes significantly more, during inference. Yet, when acceleration techniques like DPM-solver lu2022dpm are employed, reducing the number of steps below five still frequently leads to performance degradation. Policy gradient methods enable end-to-end optimization of the diffusion process, potentially approximating the target Boltzmann distribution using fewer denoising steps. However, the lack of strong intermediate supervision can lead to local optima and mode collapse.
@@
In this work, we present DACER2: a significant step toward achieving stronger performance with fewer diffusion steps. The key contributions of this paper are the following: 1) We propose a Q-gradient field objective as an auxiliary optimization target to enhance diffusion policy training at each diffusion time step. 2) We propose a time-weight function $w(t)$ that takes the current diffusion time step as input to alleviate the problem that the Q-gradient field is independent of the diffusion time. This approach aligns with the denoising process requirements: higher amplitudes during early stages and precise control for smaller amplitudes in later stages. 3) We evaluated the efficiency and generality of our method on the popular MuJoCo benchmark. Compared with DACER wang2024diffusion , QVPO ding2024diffusion , QSM psenka2023learning , DIPO yang2023diffusion , DSAC duan2023dsac , TD3 fujimoto2018addressing , PPO schulman2017proximal , and SAC haarnoja2018soft , our approach achieved state-of-the-art (SOTA) performance. 4) We evaluated the training and inference times of all diffusion-based algorithms under identical hardware configurations using the PyTorch framework. Our method achieved 60.6% faster inference and 41.7% faster training compared to DACER, while also demonstrating the fastest inference efficiency. Detailed results are presented in Fig. 4.
@@ 2.1 Reinforcement Learning with Soft Policy
RL problems are commonly modeled as Markov Decision Processes (MDPs) sutton2018reinforcement . An infinite-horizon MDP is defined by a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ is the state space and $\mathcal{A}$ is the action space, both assumed bounded and potentially continuous. $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ denotes the transition dynamics, specifying the probability distribution $P(\cdot\mid s_{t},a_{t})$ over next states, with $\Delta(\mathcal{S})$ representing the set of distributions over $\mathcal{S}$. $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, and $\gamma\in[0,1)$ is the discount factor.
The objective in RL is to find a policy $\pi$ that maximizes the expected discounted cumulative reward:
$J_{\pi}=\mathbb{E}_{(s_{i\geq t},a_{i\geq t})\sim\pi}\Big{[}\sum_{i=t}^{\infty}\gamma^{i-t}r(s_{i},a_{i})\Big{]},$ (1)
The state-action value function, or Q-function, for a policy $\pi$ is defined as:
$Q(s,a)=\mathbb{E}_{\pi}\Big{[}\sum_{i=0}^{\infty}\gamma^{i}r(s_{i},a_{i})\mid s_{0}=s,a_{0}=a\Big{]}.$ (2)
A central challenge in online RL is balancing exploration and exploitation to optimize long-term performance. One compelling strategy involves learning a policy that aims to approximate a soft policy haarnoja2017reinforcement ; haarnoja2018soft ; duan2021distributional ; duan2023dsac ; psenka2023learning ; ma2025soft ; messaoud2024s . Such target soft policies are typically formulated as a Boltzmann distribution, where the desired policy distribution is proportional to the exponentiated state-action value function:
@@
In the context of maximum entropy RL, this $Q(s,a)$ is often extended to a soft Q-function, denoted $Q_{\text{soft}}(s,a)$, which incorporates an entropy bonus into the standard value function to explicitly encourage exploration. The temperature parameter $\alpha>0$ controls the stochasticity of this target soft policy. The learned policy, $\pi$, is then trained to match this implicitly defined $\pi_{\text{soft}}$.
@@ 2.2 Diffusion Models as Expressive Policy
Diffusion models ho2020denoising ; song2020score ; wang2024diffusion conceptualize data generation as a stochastic process where data samples are iteratively reconstructed via a parameterized reverse-time stochastic differential equation (SDE). Although both forward and reverse diffusion processes are theoretically integral to these models, recent work by He et al. chen2024deconstructing highlights that their expressive power primarily stems from the reverse-time denoising dynamics, rather than the forward-time noising process. Accordingly, our analysis and modeling efforts concentrate on the reverse diffusion process.
@@
$d\boldsymbol{x}=\left[f(\boldsymbol{x},t)-g(t)^{2}\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})\right]dt+g(t)\,d\omega(t),$ (4)
where $f(\boldsymbol{x},t)$ represents the drift term, $g(t)$ denotes the time-dependent diffusion coefficient, $\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$ is the score function, and $d\omega(t)$ is the standard Wiener process. The term $\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$, also known as the score function, plays a central role in guiding the reverse diffusion dynamics. It is important to note that this equation represents the general form of the reverse-time SDE; the specific construction of terms such as $f(\boldsymbol{x},t)$ and $g(t)$ can vary across different diffusion model algorithms.
@@
$d\boldsymbol{a}_{t}=\left[f(\boldsymbol{a}_{t},t)-g(t)^{2}S_{\theta}(\boldsymbol{s},\boldsymbol{a}_{t},t)\right]dt+g(t)\,d\omega(t),$ (5)
@@
where $S_{\theta}(\boldsymbol{s},\boldsymbol{a}_{t},t)$ is a neural network designed to approximate the gradient $\nabla_{\boldsymbol{a}_{t}}\log p_{t}(\boldsymbol{a}_{t}|\boldsymbol{s})$. Actions can be sampled from the diffusion policy $\pi_{\theta}(\boldsymbol{a}_{0}|\boldsymbol{s})$ by solving the following integral:

### v2

@@ Abstract
Due to their expressive capacity, diffusion models have shown great promise in offline RL and imitation learning. Diffusion Actor-Critic with Entropy Regulator (DACER) extended this capability to online RL by using the reverse diffusion process as a policy approximator, achieving state-of-the-art performance. However, it still suffers from a core trade-off: more diffusion steps ensure high performance but reduce efficiency, while fewer steps degrade performance. This remains a major bottleneck for deploying diffusion policies in real-time online RL. To mitigate this, we propose DACERv2, which leverages a Q-gradient field objective with respect to action as an auxiliary optimization target to guide the denoising process at each diffusion step, thereby introducing intermediate supervisory signals that enhance the efficiency of single-step diffusion. Additionally, we observe that the independence of the Q-gradient field from the diffusion time step is inconsistent with the characteristics of the diffusion process. To address this issue, a temporal weighting mechanism is introduced, allowing the model to effectively eliminate large-scale noise during the early stages and refine its outputs in the later stages. Experimental results on OpenAI Gym benchmarks and multimodal tasks demonstrate that, compared with classical and diffusion-based online RL algorithms, DACERv2 achieves higher performance in most complex control environments with only five diffusion steps and shows greater multimodality.
@@ 1 Introduction
Figure 1: Efficiency and Performance. The horizontal axis represents the training or inference time (increasing from right to left), while the vertical axis shows the normalized Total Average Return (TAR). The training time is the per-step computational cost on OpenAI Gym tasks, excluding the time spent on environment interaction. The inference time is measured as the latency required for the policy network to output an action given a single state as input. DACERv2 achieve outstanding performance.
@@
Energy-based models are well-suited as agent policy functions due to their powerful representational capailities. Learning a policy to approximate the corresponding energy-based target distribution allows for modeling complex and multimodal action patterns without relying on restrictive parametric assumptions, especially in continuous action spaces. This enhanced expressiveness can significantly improve exploration by enabling the agent to discover and leverage diverse behavioral strategies. However, effectively approximating such an expressive soft policy presents notable challenges. A key difficulty lies in how to efficiently and accurately sample from the target distribution. While algorithms such as Soft Actor-Critic (SAC) (Haarnoja et al., 2018) and Distributional Soft Actor-Critic (DSAC) (Duan et al., 2021; 2025) aim to approximate the soft-target distribution, they typically represent the policy as a simple Gaussian, enabling analytical entropy computation. This choice is computationally efficient but falls short in modeling complex and multimodal behavior. Meanwhile, due to their strong representational capacity, diffusion models have emerged as a promising policy class for continuous control, commonly referred to as diffusion policies (Ren et al., 2024; Li et al., 2024; Lu et al., 2025).
@@
Existing methods for training diffusion policies can be broadly categorized into two groups: score-matching and end-to-end policy gradient approaches. In the first group, QVPO (Ding et al., 2024) proposes using Q-weighted imitation learning samples to improve policy learning. QSM (Psenka et al., 2023) directly aligns the score functions with the gradients of the learned Q-functions and uses Langevin dynamics for sampling. DIPO (Yang et al., 2023a) updates the replay buffer using action gradients and improves the performance of the policy through a diffusion loss. In the second group, DACER (Wang et al., 2024) directly backward the gradient through the reverse diffusion process and proposes a Gaussian mixture model (GMM) entropy regulator to balance exploration and exploitation. DIME (Celik et al., 2025) derives an approximate maximum-entropy lower bound, directly integrating the maximum-entropy RL framework with the diffusion policy. However, diffusion policies typically require a large number of diffusion steps to maintain strong performance, which results in low inference efficiency. Although acceleration techniques such as DPM-Solver (Lu et al., 2022) can reduce the number of diffusion steps, this often comes at the cost of performance degradation. As a result, previous methods struggle to escape the dilemma between performance and time-efficiency.
@@
To tackle this issue, we present DACERv2, a highly efficient diffusion-based RL algorithm that achieves comparable or superior performance with only a few diffusion steps, as shown in Fig. 1. The key contributions of this paper are the following: 1) We propose a Q-gradient field objective as an extra intermediate supervisory signal to enhance the efficiency of single-step diffusion. 2) Since the Q-gradient field is independent of the diffusion time, we propose a temporal weighting mechanism that takes the current diffusion time step as input. This mechanism aligns with the requirements of the diffusion denoising process: higher denoising amplitudes during early stages and lower denoising amplitudes for precise control in later stages. 3) We evaluate the performance of our method on the OpenAI Gym benchmark (Brockman, 2016). Compared with both diffusion-based and classical algorithms like DACER (Wang et al., 2024), QVPO (Ding et al., 2024), DIME (Celik et al., 2025), QSM (Psenka et al., 2023), DIPO (Yang et al., 2023a), DSAC (Duan et al., 2025), PPO (Schulman et al., 2017), and SAC (Haarnoja et al., 2018), our approach achieved state-of-the-art (SOTA) performance in most complex control tasks. 4) We evaluate the training and inference times of all diffusion-based algorithms under identical hardware configurations using the PyTorch framework. While achieving stronger overall performance, our method reduces training time by 47.0% and 41.7%, and inference time by 55.0% and 60.6%, compared with DIME and DACER, respectively.
@@ 2.1 Reinforcement Learning with Soft Policy
RL problems are commonly modeled as Markov Decision Processes (MDPs) (Sutton & Barto, 2018; Li, 2023). An infinite-horizon MDP is defined by a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ is the state space and $\mathcal{A}$ is the action space, both assumed bounded and potentially continuous. $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ denotes the transition dynamics, specifying the probability distribution $P(\cdot\mid s_{t},a_{t})$ over next states, with $\Delta(\mathcal{S})$ representing the set of distributions over $\mathcal{S}$. $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. The behavior of agent is characterized by a policy $\pi:\mathcal{S}\mapsto A$, which defines the process of selecting action $a$ given the state $s$.
To evaluate the value of taking an action $a$ in a given state $s$ under policy $\pi$, the action-value function $Q^{\pi}(s,a)$ is introduced, which represents the expected cumulative discounted reward, defined as follows:
$Q^{\pi}(s,a)=\mathbb{E}_{\pi}\Big[\sum_{i=0}^{\infty}\gamma^{i}r(s_{i},a_{i})\mid s_{0}=s,a_{0}=a\Big].$ (1)
A key challenge in online RL is the trade-off between exploration, gathering information for future gains, and exploitation, maximizing returns based on current knowledge. One compelling strategy involves learning a policy that aims to approximate a soft policy (Haarnoja et al., 2017; 2018; Ma et al., 2025; Messaoud et al., 2024). Such target soft policies are typically formulated as a Boltzmann distribution, where the desired policy distribution is proportional to the exponentiated state-action value function:
@@
The target of soft policy is to minimize the per-state KL divergence $D_{\mathrm{KL}}\!\left(\pi(\cdot|s)\,\|\,\frac{\exp(Q(s,\cdot)/\alpha)}{Z(s)}\right)$, where $Z(s)$ is the normalization coefficient. This KL-divergence minimization problem is equivalent to maximizing a final objective function that balances value maximization and entropy regularization:
$J(\pi)=\mathbb{E}_{(s,a)\sim\pi}\big[Q(s,a)\big]+\alpha\cdot\mathcal{H}(\pi(\cdot|s)).$ (3)
Diffusion policies are able to model complex policy distributions, but their entropy is analytically intractable. Fortunately, in methods like DACER (Wang et al., 2024), maximizing the $Q$-value objective under specific entropy regularization likewise produces a Boltzmann policy. See Theorem 1 for further theoretical details.
@@ 2.2 Diffusion Models as Expressive Policy
Diffusion models (Ho et al., 2020; Song et al., 2020; Wang et al., 2024) conceptualize data generation as a stochastic process where data samples are iteratively reconstructed via a parameterized reverse-time stochastic differential equation (SDE). Although both forward and reverse diffusion processes are theoretically integral to these models, recent work (Chen et al., 2024) highlights that their expressive power primarily stems from the reverse-time denoising dynamics, rather than the forward-time noising process. Accordingly, our analysis and modeling efforts concentrate on the reverse diffusion process.
@@
$d\bm{x}=\left[f(\bm{x},t)-g(t)^{2}\nabla_{\bm{x}}\log p_{t}(\bm{x})\right]dt+g(t)\,d\omega(t),$ (4)
where $f(\bm{x},t)$ represents the drift term, $g(t)$ denotes the time-dependent diffusion coefficient, $\nabla_{\bm{x}}\log p_{t}(\bm{x})$ is the score function, and $d\omega(t)$ is the standard Wiener process. The term $\nabla_{\bm{x}}\log p_{t}(\bm{x})$, also known as the score function, plays a central role in guiding the reverse diffusion dynamics. It is important to note that this equation represents the general form of the reverse-time SDE; the specific construction of terms such as $f(\bm{x},t)$ and $g(t)$ can vary across different diffusion model algorithms.
@@
$d\bm{a}_{t}=\left[f(\bm{a}_{t},t)-g(t)^{2}S_{\theta}(\bm{s},\bm{a}_{t},t)\right]dt+g(t)\,d\omega(t),$ (5)
@@
where $S_{\theta}(\bm{s},\bm{a}_{t},t)$ is a neural network designed to approximate the gradient $\nabla_{\bm{a}_{t}}\log p_{t}(\bm{a}_{t}|\bm{s})$. Actions can be sampled from the diffusion policy $\pi_{\theta}(\bm{a}_{0}|\bm{s})$ by solving the following integral:
$\bm{a}_{0}=\bm{a}_{T}+\int_{0}^{T}\!\left[f\!\left(\bm{a}_{\tau},\tau\right)-g(\tau)^{2}\,S_{\theta}\!\left(\bm{s},\mathbf{a}_{\tau},\tau\right)\right]d\tau+\int_{0}^{T}g(\tau)\,d\omega(\tau),$ (6)
where $\bm{a}_{T}$ follows a standard normal distribution $\mathcal{N}(0,\bm{I})$.
2.3 Langevin Dynamics
Langevin dynamics represents a powerful computational framework for simulating particle motion under the joint influence of deterministic forces and stochastic fluctuations. When coupled with stochastic gradient descent, this approach gives rise to stochastic gradient Langevin dynamics (SGLD) (Welling & Teh, 2011) - an efficient sampling algorithm that leverages log-probability gradients $\nabla_{\bm{x}}\log p(\bm{x})$ to draw samples from probability distributions $p(x)$ through an iterative Markov chain process:
$\bm{x}_{t-1}=\bm{x}_{t}+\frac{\delta_{t}}{2}\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})+\sqrt{\delta}_{t}\bm{\epsilon},$ (7)
where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, $\delta_{t}$ is the step size. When $t$ range from infinity to one, $\delta_{t}\to 0$, $x_{0}$ equals to the true probability density $p(x)$.
3 Method
In this section, we explain how our method approximates the target policy distribution with fewer diffusion steps. First, we show that $\nabla_{\bm{a}_{t}}Q(\bm{s},\bm{a}_{t})$, derived from Langevin dynamics, can be incorporated into the unified SDE-based framework for action generation, thereby improving the efficiency of single-step diffusion. However, when this extra objective function is incorporated, the diffusion policy only exhibits suboptimal performance. This limitation arises because $\nabla_{\bm{a}_{t}}Q(\bm{s},\bm{a}_{t})$ remains independent of the diffusion step, whereas the score function is not. Therefore, we introduce a time-weighted mechanism to better align with the requirements of the diffusion denoising process. Lastly, we propose a practical algorithm for optimizing diffusion models.
3.1 Q-Gradient Field Guided Denoising
Using the only reverse process, the objective function of DACER is to maximize the Q-value, representing an end-to-end optimization approach without direct supervision in the intermediate diffusion steps. However, within this optimization scheme, the guidance signals at intermediate steps are implicit, as they are derived solely from the final Q-value through back-propagation, which in turn necessitates more diffusion steps to produce higher-quality control actions. To address this issue, we propose the Q-gradient field function as an extra training loss to enhance the efficiency of single-step diffusion. At the end of Section 2.1, we explain why, when the global policy entropy is fixed, the optimal policy for maximizing the Q-value theoretically follows a Boltzmann distribution. Importantly, this conclusion holds for policy families of arbitrary forms and naturally suits the SDE-based policy families.
From another perspective, Langevin dynamics can be regarded as a special form of an SDE-based policy, providing an efficient method for sampling actions from Boltzmann distributions (Hinton, 2002):
$\pi(\bm{a}|\bm{s})=\frac{e^{\frac{1}{\alpha}Q(\bm{s},\bm{a})}}{Z(\bm{s})},$ (8)
where $\alpha>0$ is the temperature coefficient, $Q(\bm{s},\bm{a})$ is the state action value function and $Z(\bm{s})$ is the partition function that normalizes the distribution. The formula derived by taking the partial derivative of both sides of Eq. (8) with respect to $\bm{a}$ can be expressed as
$\nabla_{\bm{a}}\log\pi(\bm{a}|\bm{s})=\frac{1}{\alpha}\nabla_{\bm{a}}Q(\bm{s},\bm{a}).$ (9)
Substituting Eq. (9) into Eq. (7), we can obtain the sampling process for $\pi(\bm{a}|\bm{s})$:
$\bm{a}_{t-1}=\bm{a}_{t}+\frac{\delta_{t}}{2\alpha}\nabla_{\bm{a}}Q(\bm{s},\bm{a})+\sqrt{\delta}_{t}\bm{\epsilon}.$ (10)
In summary, Langevin dynamics can be regarded as a particular solution within the family of SDE-based policies. This connection motivates the use of $\nabla_{\bm{a}}Q(\bm{s},\bm{a})$ as an extra learning objective to guide the training of SDE-based policies, thereby introducing additional supervision signals into the intermediate diffusion step. Consequently, the efficiency of single-step diffusion can be improved, enabling comparable or even superior performance to previous algorithms with fewer diffusion steps.
Moreover, in highly unstable environments that exhibit extreme sensitivity to minor action perturbations, the Q-gradient estimation can become volatile, potentially hindering the algorithm’s convergence to the optimal policy (Ding et al., 2024; Ma et al., 2025). When the diffusion process is restricted to only a few steps, a policy trained solely on the Q-gradient often struggles to converge. For these reasons, we adopt it only as an auxiliary guidance in policy training.
3.2 Time-weighted mechanism
In the previous subsection, we propose incorporating the gradient term $\nabla_{\bm{a}_{t}}Q(\bm{s},\bm{a}_{t})$ as an auxiliary objective when training the SDE-based policy. However, experimental results show that directly employing this objective yields suboptimal performance, as shown in Fig. 4(b). We attribute this to the Q-gradient being independent of the diffusion time step, whereas the score function is not. Such time invariance fails to satisfy the varying denoising requirements across the diffusion process. Specifically, in the later stages of diffusion process, the denoising intensity should naturally decrease as the action distribution approaches the optimal policy.
To address this issue, we introduce a time-weighted mechanism that modulates the influence of Q-gradient guidance based on the diffusion time step, allowing for more precise control over the denoising process. Inspired by the design approach for the step size $\delta_{t}$ in Eq. (7), we can similarly design our time-weighted mechanism using the commonly employed exponential decay function (Welling & Teh, 2011; Teh et al., 2016):
$w(t)=\exp(c\cdot t+d),$ (11)
where $t$ denotes the current diffusion step. The hyperparameters $c$ and $d$ are chosen inspired by the variance-preserving beta schedule in DDPM (Ho et al., 2020) and depend only on the number of diffusion steps. The specific setting is presented in Appendix D.
Furthermore, to improve the stability of the training process, we normalize $\nabla_{\bm{a}_{t}}Q(\bm{s},\bm{a}_{t})$ by its norm: